# A Quantum Approximate Optimization Method For Finding Hadamard Matrices

### v1

@@ 3 Experiments
We do the experiments using both of simulators and quantum hardware. In the latest case, we implement simple experiment on low qubit number of IBM quantum computer. Before implementing the quantum circuit in Hadamard search, which involves several k-body interaction terms, we have tested its elementary circuits displayed in Fig.1 individually. The performance for each of these circuits are as expected, i.e., the distribution of solutions shows the validity of the circuit. Detail of the results can be found in the Supplementary Information,
The results of the algorithms discussed in this paper are in the form of $L$-length bit strings, so that there are $2^{L}$ possible combination of them. A result is labelled as valid or correct when the value of the non-negative error/energy function related to the algorithm, i.e. the value of the Williamson or Turyn cost function, is equal to zero. Otherwise, it is labelled as wrong. To measure the performance of the algorithm in producing correct solutions, we can compare it to an algorithm that randomly generate every possible combination of an $L$ length bit string. Therefore, for a given x-algorithm to measure, we propose xRAR (x-Algorithm to Random-Algorithm Ratio) as a performance metric.
@@
The random algorithm generates $2^{L}$ number of bit strings and we can evaluate whether each of the bit string is valid or not. When we know beforehand that there are $S_{R}$ number of valid solutions among the $2^{L}$ random bit strings, assuming a uniform probability distribution of the bit strings, the probability $P_{R}$ to find a valid solution is $P_{R}=\frac{S_{R}}{2^{L}}$. On the other hand, the solutions generated by a quantum circuit that represents the x-algorithm are also probabilistic, and we can calculate the probability $P_{x}$ of valid solution of the x-algorithm. Therefore, the value of xRAR; which conceptually is illustrated by Fig.2 (a), is given as follows
@@
The value of xRAR in Eq.(16) is a positive real number, where $0<\textit{xRAR}<1$ indicates that the x- algorithm is worse than the random algorithm $R$, $\textit{xRAR}=1$ it has a similar performance to the random algorithm, and $\textit{xRAR}>1$ means that the x-algorithm is better than random algorithm $R$.
The execution steps of the QAOA in the experiments is displayed in Fig.2 (b). When running the quantum algorithm, either in a simulator or in a real quantum computer, we repeat the run $N$ times, which is called as the number of shots. It yields $N$ number of solutions/ bit strings and its corresponding energy/error. A particular block/part of the algorithm will calculate the average or expectation, where we also can check whether particular bit strings or solution achieve a minimum energy shown as the block Extraction of Measured Qubits in the figure. We can count the number of the valid solution in each iteration step and after the lowest average energy for a particular experiment setup is achieved. The number of correct solutions can be used for determining the performance of the algorithm.
In addition to XRAR, we also evaluate the performance of the algorithm with an error metric. This error metric is accumulated or total objective values of all generated solutions, with the objective measure are given by either the Williamson or Turyn cost functions given by Eq.(3) and Eq.(6), respectively. The value is normalized by the maximum of each cost function, then compared to the value of random algorithm which is calculated by exhaustive search. Therefore, for a particular order of H-SEARCH that generates $N_{Q}$-length bit string solution with maximum objective error of $E_{max}$ and total error for all possible $2^{N_{Q}}$-bit string equal to $E_{tot}$, we have normalization factor $E_{max}2^{N_{Q}}$, the average error of $E_{tot}/2^{N_{Q}}$, and normalized average error of $E_{tot}/(2^{N_{Q}}E_{max})$. The normalized average error is used to compare the performance of algorithms with various number of shots (sampling).
@@
An exhaustive search to all possible $2^{8}$ bit strings that yields minimum energy; and therefore correct bit strings, found $64$ WBH-sequences as valid solutions. This result yields the probability value to find the solution of 8-length uniformly distributed bit strings $P_{R}=\frac{1}{4}$; therefore, the maximum QRAR performance is $\frac{1}{P_{R}}=4$. It is also found that the maximum value of the cost function is $18$ and the total error of $1024$; therefore, the normalized average error is equal to $0.2222$. The Hamiltonian of this problem is given by
@@
Note that in the minimization, the constant term can be dropped without affecting the result. We will do some experiments for this case with both of the simulator and real quantum computer.
@@
We also implemented the algorithm of finding 12-order Williamson matrix in a quantum computer hardware. An IBM quantum computer, in this case is the IBM-Brisbane machine powered by a 127 qubits Eagle r.3 of version 1.1.6 quantum processor, was employed. The processor’s qubits mean coherence time are $T_{1}\approx 227\mu s,T_{2}\approx 130\mu s$ with median ECR error $\approx 7\times 10^{-3}$ and median SX error $\approx 2\times 10^{-4}$. We also repeat the run 10 times and the number of shots in the hardware is set to 1024. The results is displayed in Fig. 4(b), which is the quantum hardware (QPU) performance for the QAOA 12-order Williamson method with number of layers $1,2,3$ and $4$. This figure shows that although the mean error of QAOA implemented on hardware (blue dotted line with ”$\times$” symbols) are consistently lower than the mean error of random algorithm (blue dotted line in the upmost part), the mean XRAR performance (red solid line with red circle symbols) is sometimes only slightly better than the random algorithm bound (red solid line) for number of layer of 1 and 3, and worse for 2 and 4. Since initialization of the angle can influence the final results, in term of XRAR, we also display the maximum XRAR for each repeated 10 times run. The max RAR performance initially higher than random but then tend to decrease when the number of layers are increased. This shows the circuit depth enlarge noise in the qubits.
(a) 12-order Williamson
(b) 36-order Baumert-Hall
Figure 5: The 12-order Williamson/36-order Baumert-Hall Hadamard Matrices found by the algorithms. Both simulation and hardware found identical H-matrices.
@@
The corresponding quantum circuit can be constructed automatically using a construction algorithm, which the one layer circuit is shown in Fig. 1 (f). We then run the QAOA on both of quantum simulator and quantum hardware, which successfully found the 44-order Turyn H-matrix. As shown in the QAOA flowchart in Fig. 2, normally we need to put the process done in the quantum computing inside the optimization, so that an exclusive access to this device is mandatory. However, since we used a public access to 5 qubits IBM quantum computer, such dedicated access was not permitted. Therefore, we only implemented the optimized quantum circuits obtained in the simulation into the 5 qubits IBM Quito.
@@
Figure 6: Results for 44-order Turyn Method
The 44-order H matrix for both of the simulator and hardware are identical, which is shown in Fig. 6: (a) implementation in IBM-Quito, and (b) obtained order-44 of the Turyn H-matrix. The histogram indicates that the number of the valid solution, i.e. the bit string ”11100”, is equal to 11; which means that it consists of about $6\%$ correct solution. Since random algorithm would have yield only $3.1\%$, the quantum processor indicates a little bit more advantage than the classical.
The PEL of the QAOA for high order H-matrices are more rough than the lower ones shown previously, implying that selection of initial parameters is both difficult and crucial. In this subsection, we show the results of finding 44-order Williamson/132-order Baumert-Hall matrices. We use only 1-layer and choose various initialization parameters. The number of required qubits to implement is 24. After setting the number of sampling to 10,000 shots, we obtained the mean xRAR on 10 different parameter initialization is equal to 1.14 with the maximum value of 3.50. One of the obtained matrix is displayed in Fig. 7, where (a) shows 44-order Williamson and (b) the 132-order Baumert-hall matrices.
@@
(b) 132-order Baumert-Hall Check Matrix
@@ 4 Discussion
We have shown a possibility to implement a Hadamard matrix searching method on a quantum computer simulator. Experiments on a quantum computer simulator shows an exponential improvement of the QAOA-based quantum algorithm over random search algorithm. At the final stage, we have run and find a 44-order Williamson or 132-order Baumert-Hall matrices successfully.

### v2

@@ 3 Experiments
We conducted experiments using both simulators and quantum hardware. In the latter case, we implemented a simple experiment on an IBM quantum computer. Before implementing the quantum circuit for the Hadamard search, which involves several $k$-body interaction terms, we tested its elementary circuits shown in Fig. 1 individually. The performance of each circuit met expectations, with the solution distributions confirming the circuit’s validity. Detailed results are provided in the Supplementary Information.
The outputs of the algorithms discussed in this paper are $L$-length bit strings, resulting in $2^{L}$ possible combinations. An output string is considered valid or correct if the value of the associated non-negative error or energy function—such as the Williamson or Turyn cost function—is zero. Otherwise, it is labeled as incorrect (wrong). To evaluate the algorithm’s performance in producing correct solutions, we compare it to an algorithm that randomly generates all possible $L$-length bit strings. Accordingly, we introduce the xRAR (x-Algorithm to Random-Algorithm Ratio) as a performance metric for a given x-algorithm.
@@
The random algorithm generates $2^{L}$ number of bit strings and we can evaluate whether each of the bit string is valid or not. If there are $S_{R}$ valid solutions among the $2^{L}$ random bit strings and we assume a uniform probability distribution, the probability of finding a valid solution, $P_{R}$, is given by $P_{R}=\frac{S_{R}}{2^{L}}$. Similarly, the solutions generated by a quantum circuit that represents the x-algorithm are also probabilistic, and we can calculate the probability $P_{x}$ of valid solution of the x-algorithm. Therefore, the value of xRAR; which conceptually is illustrated by Fig.2 (a), is given as follows
@@
The value of xRAR in Eq. (16) is a positive real number. An xRAR value where $0<\textit{xRAR}<1$ indicates that the x-algorithm performs worse than the random algorithm $R$, $\textit{xRAR}=1$ signifies comparable performance to the random algorithm, and $\textit{xRAR}>1$ suggests that the x-algorithm outperforms the random algorithm $R$.
The execution steps of the QAOA used in the experiments are shown in Fig. 2 (b). When running the quantum algorithm, either on a simulator or a real quantum computer, we repeat the process $N$ times, referred to as the number of shots. This produces $N$ solutions or bit strings, each with a corresponding energy or error. A specific part of the algorithm computes the average or expectation value, allowing us to determine whether any of the bit strings achieve the minimum energy, as indicated in the Extraction of Measured Qubits block in the figure. We can count the number of valid solutions at each iteration step and also after reaching the lowest average energy for a given experimental setup. The number of correct solutions is then used to evaluate the algorithm’s performance.
In addition to xRAR, we also evaluate the performance of the algorithm using an error metric. This error metric is defined as the accumulated or total objective values of all generated solutions, with the objectives measured by either the Williamson or Turyn cost functions, as given by Eq. (3) and Eq. (6), respectively. The error is normalized by the maximum value of each cost function and then compared to the value obtained by a random algorithm through exhaustive search.
For a particular order of H-SEARCH that generates $N_{Q}$-length bit string solutions with a maximum objective error of $E_{max}$ and a total error for all possible $2^{N_{Q}}$ bit strings equal to $E_{tot}$, the normalization factor is $2^{N_{Q}}E_{max}$. The average error is given by $E_{tot}/2^{N_{Q}}$, and the normalized average error is $E_{tot}/(2^{N_{Q}}E_{max})$. This normalized average error is used to compare the performance of algorithms with varying numbers of shots (samples).
@@
An exhaustive search to all possible $2^{8}$ bit strings that yields minimum energy; and therefore correct bit strings, found $64$ WBH-sequences as valid solutions. This result yields the probability value to find the solution of 8-length uniformly distributed bit strings $P_{R}=\frac{1}{4}$; therefore, the maximum QRAR performance is $\frac{1}{P_{R}}=4$. It is also found that the maximum value of the cost function is $18$ and the total error of $1,024$; implying that the normalized average error is equal to $0.2222$. The Hamiltonian of this problem is given by
@@
Note that in the minimization, the constant term can be dropped without affecting the result. We will perform some experiments for this case with both of the simulator and the real quantum computer.
@@
We also implemented the algorithm of finding 12-order Williamson matrix in a quantum computer hardware. An IBM quantum computer, in this case is the IBM-Brisbane machine powered by a 127 qubits Eagle r.3 of version 1.1.6 quantum processor, was employed. The processor’s qubits mean coherence time are $T_{1}\approx 227\mu s,T_{2}\approx 130\mu s$ with median ECR error $\approx 7\times 10^{-3}$ and median SX error $\approx 2\times 10^{-4}$. We also repeat the run 10 times and the number of shots in the hardware is set to 1024. The results is displayed in Fig. 4(b), which is the quantum hardware (QPU) performance for the QAOA 12-order Williamson method with number of layers $1,2,3$ and $4$. This figure shows that although the mean error of QAOA implemented on hardware (blue dotted line with ”$\times$” symbols) are consistently lower than the mean error of random algorithm (blue dotted line in the upmost part), the mean XRAR performance (red solid line with red circle symbols) is sometimes only slightly better than the random algorithm bound (red solid line) for number of layer of 1 and 3, and worse for 2 and 4. Since initialization of the angle can influence the final results, in term of XRAR, we also display the maximum XRAR for each repeated 10 times run. The max RAR performance initially higher than random but then tend to decrease when the number of layers are increased. This shows the circuit depth increases the noise level of the qubits.
(a) 12-order Williamson H-matrix
(b) 36-order Baumert-Hall H-matrix
Figure 5: The 12-order Williamson and 36-order Baumert-Hall Hadamard Matrices found by the proposed algorithms. Both simulation and hardware found identical H-matrices. In the figure, white boxes represent ”+1” elements and the black ones represent ”-1”.
@@
The corresponding quantum circuit can be automatically constructed using a construction algorithm, with a single-layer circuit depicted in Fig. 1 (f). We then ran the QAOA on both a quantum simulator and quantum hardware, successfully identifying the 44th-order Turyn H-matrix. As shown in the QAOA flowchart in Fig. 2, normally we need to put the process done in the quantum computing inside the optimization, requiring exclusive access to the device. However, since we used a public access to a 5 qubits IBM quantum computer, such dedicated access was not permitted. Therefore, we only implemented the optimized quantum circuits obtained in the simulation into the 5 qubits IBM Quito.
@@
Figure 6: Results for 44-order Turyn method: (a) Histogram of the output when running the algorithm in IBM Quito. The inset shows qubit configuration on the quantum device. (b) A Turyn Hadamard matrix of order-44 found by the proposed method running on IBM Quito
The obtained 44-order H matrix for both of the simulator and hardware are identical, which is shown in Fig. 6: (a) output histogram of implemented algorithm in IBM-Quito, and (b) result of order-44 of the Turyn H-matrix. The histogram in Fig. 6 (a) indicates that the number of the valid solution, i.e. the bit string ”11100”, is equal to 11; which means that it consists of about $6\%$ valid solution. Since random algorithm would have yield only $3.1\%$, the quantum processor indicates a slight advantage over the classical one. The PEL of the QAOA for higher-order H-matrices is more irregular compared to that for lower-order matrices, as shown previously. This suggests that selecting initial parameters is both challenging and crucial. In the followings, we show the results of finding 44-order Williamson/132-order Baumert-Hall matrices using a simulator. We used a single-layer QAOA and experimented with various initial parameters, selecting the best-performing configuration. The required number of qubits to implement this scheme is 24. After setting the number of sampling to 10,000 shots, we obtained the mean xRAR on 10 different parameter initialization is equal to 1.14 with the maximum value of 3.50. One of the obtained matrix is displayed in Fig. 7, where (a) shows 44-order Williamson and (b) the 132-order Baumert-hall matrices.
@@
(b) 132-order Baumert-Hall Hadamard matrix
@@ 4 Discussion
We have demonstrated the feasibility of implementing Hadamard matrix search algorithms on a circuit-based quantum computer, both in a simulator and on actual quantum hardware. Within the framework of quantum optimization, we utilized the Quantum Approximate Optimization Algorithm (QAOA) to construct the Hamiltonian for optimization, as previously described in our work [24]. This Hamiltonian was then implemented in quantum circuits and executed on circuit-based quantum computers. Due to hardware limitations and the current state of noisy qubits, the quantum hardware was only tested at the lowest order. However, the quantum simulator was able to successfully execute higher-order cases
Experimental results suggest that the difficulty in finding H-matrices using QAOA algorithms arises from the non-smoothness of the energy landscape (PEL), which becomes increasingly pronounced in higher-order cases. While the Turyn-based method is more qubit-efficient than the Williamson and Baumert-Hall (WBH) method, its more irregular energy landscape makes finding Turyn’s solution more challenging.
Experiments with the lowest-order WBH case (as shown in Fig. 4) on the quantum simulator demonstrate that increasing the number of layers consistently enhances performance. This improvement is indicated by the xRAR metric, which asymptotically approaches the performance limit as the number of layers increases. However, implementing the algorithm on a real quantum device did not replicate this performance. With a single layer, the algorithm performed slightly better than a random algorithm on average, but this advantage diminished as the number of layers increased—unless only the best performance from multiple iterations was selected. However, the advantage of using more than one layer also disappears. These results suggest that increasing the number of layers on a NISQ device provides only limited benefits.
Late last year, in 2023, IBM successfully built a 1,121-qubit processor, known as the Condor quantum processor, although the issue of noise remains unresolved. More recently, quantum error correction experiments have reached the threshold for the surface code [30]. If these trends continue, it is likely that some of the currently unknown Hadamard matrices will eventually be discovered. The QAOA method would require 336 qubits to find the lowest unknown 668-order H-matrix using the Williamson method, or 157 qubits using the Turyn method. In terms of qubit numbers, this is within the reach of current technology. However, noise remains a significant obstacle to implementation. Nonetheless, it would be exciting to explore this domain further, particularly with exclusive access to a quantum device capable of running QAOA at full scale.